OpenStudy (anonymous):

Two boats leave port at the same time, one hitting north at 32 knots ( nautical miles per hour), the other south at 45 knots. How long will it take them to be 770 nautical miles apart?

OpenStudy (anonymous):

10 hours

OpenStudy (anonymous):

is it with 45 +32 divided byt 770?

OpenStudy (anonymous):

see distance = velocity x time // here relative velocity = 32 + 45 => 770 = (32+45)t => t = 770/77 = 10hours